OpenStudy (anonymous):

Fine the volume of thefigure, it is a rectangler and the measrements are 3x^2, 4x^3, 5x^5

OpenStudy (anonymous):

The volume is length * width * height. So all you do is multiply the three dimensions. \[3x^2 \times 4x^3 \times 5x^5 = 60x^{10}\] As you can see, the volume will depend on x.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

No problem. Best response? :)